OpenStudy (anonymous):

3. Find the mean of the following numbers. 5, 11, 7, 12, 1, 0 3. Find the mean of the following numbers. 5, 11, 7, 12, 1, 0 @Mathematics

OpenStudy (anonymous):

add them then divide the answer by their number

OpenStudy (ashish56):

the mean =(5+11+7+12+1+0)/6=36/6=6

OpenStudy (anonymous):

thanks

OpenStudy (ashish56):

mp